# The Nonlinear Filter Model of Stream Cipher Redivivus

### v1

@@ Abstract
The nonlinear filter model is an old and well understood approach to the design of secure stream ciphers. Extensive research over several decades has shown how to attack stream ciphers based on this model and has identified the security properties required of the Boolean function used as the filtering function to resist such attacks. This led to the problem of constructing Boolean functions which provide adequate security and at the same time are efficient to implement. Unfortunately, over the last two decades no good solutions to this problem appeared in the literature. The lack of good solutions has effectively led to nonlinear filter model becoming more or less obsolete. This is a big loss to the cryptographic design toolkit, since the great advantages of the nonlinear filter model are its simplicity, well understood security and the potential to provide low cost solutions for hardware oriented stream ciphers. In this paper we construct balanced functions on an odd number $n\geq 5$ of variables with the following provable properties: linear bias equal to $2^{-\lfloor n/2\rfloor-1}$, algebraic degree equal to $2^{\lfloor\log_{2}\lfloor n/2\rfloor\rfloor}$, algebraic immunity at least $\lceil(n-1)/4\rceil$, fast algebraic immunity at least $1+\lceil(n-1)/4\rceil$, and can be implemented using $O(n)$ NAND gates. The functions are obtained from a simple modification of the well known class of Maiorana-McFarland bent functions. By appropriately choosing $n$ and the length $L$ of the linear feedback shift register, we show that it is possible to obtain examples of stream ciphers which are $\kappa$-bit secure against known types of attacks for various values of $\kappa$. We provide concrete proposals for $\kappa=80,128,160,192,224$ and $256$. For the $80$-bit, $128$-bit, and the $256$-bit security levels, the circuits for the corresponding stream ciphers require about 1743.5, 2771.5, and 5607.5 NAND gates respectively. For the $80$-bit and the $128$-bit security levels, the gate count estimates compare quite well to the famous ciphers Trivium and Grain-128a respectively, while for the $256$-bit security level, we do not know of any other stream cipher design which has such a low gate count.
@@ 1 Introduction
Extensive research carried out over the last few decades has shown several approaches to cryptanalysing the nonlinear filter model of stream ciphers. The initial line of attack was based upon determining the linear complexity of the produced keystream. For an LFSR of length $L$ and a Boolean function of algebraic degree $d$, under certain reasonable and easy-to-ensure conditions, the linear complexity of the keystream was shown to be at least (${L\choose d}$ (see [37]). Using large enough values of $L$ and $d$, linear complexity based attacks can be made infeasible. The second phase of attacks consisted of various kinds of (fast) correlation attacks. Starting with the first such attack in 1985 [41], a long line of papers [35, 23, 17, 28, 29, 15, 12, 30, 11, 16, 42, 45, 44, 31, 32] explored various avenues for mounting correlation attacks. Surveys of some of the older attacks appear in [9, 10, 33, 2]. The resistance to correlation attacks is mainly determined by the linear bias $\varepsilon$ of the Boolean function $f$. The linear bias is determined by the nonlinearity of the function $f$; the higher the nonlinearity, the lower the linear bias. The third phase of attacks started in 2003 with the publication of the algebraic attack [19] and was soon followed by the publication of the fast algebraic attack [18]. Resistance to these attacks required the function $f$ to possess high (fast) algebraic immunity.
The various attacks mentioned above have posed the following design challenge for a Boolean function to be used in the nonlinear filter model of stream ciphers. Construct balanced Boolean functions which achieve a good combination of high nonlinearity and high algebraic resistance and are also very efficient to implement. Unfortunately, since the time algebraic attacks were proposed in the early-2000s, no good solutions to the design challenge for Boolean functions have appeared in the literature. A consequence of not being able to find good solutions to the design challenge is that the nonlinear filter model of stream ciphers became obsolete. This is somewhat unfortunate since the model is very old, well studied and its security is well understood.
We note that even though the nonlinear filter model became obsolete, use of LFSRs in the design of stream cipher has continued for both hardware and software oriented proposals (see for example [26, 4, 22]). Instead of using a Boolean function, such designs typically use a nonlinear finite state machine to filter the output of the LFSR. Some ciphers such as [8, 3] have gone further and replaced the LFSR with a nonlinear feedback shift register (NFSR).
@@
The well known Maiorana-McFarland class of bent functions is defined as follows. For $m\geq 1$, let $\mathbf{X}$ and $\mathbf{Y}$ be two vectors of $m$ variables. Then a $2m$-variable Maiorana-McFarland bent function is defined to be $\langle\pi(\mathbf{X}),\mathbf{Y}\rangle\oplus h(\mathbf{X})$, where $\pi$ is a bijection from $m$-bit strings to $m$-bit strings and $h$ is any $m$-variable Boolean function. It is well-known that the nonlinearity of Maiorana-McFarland functions does not depend on the choice of $h$ nor on that of the permutation $\pi$. Bad choices of $\pi$ and $h$ (for example, $\pi$ to be the identity permutation and $h$ to be a constant function) provide functions whose algebraic immunity is low. On the bright side, we observe that it may be possible to improve the algebraic resistance by properly choosing $\pi$ and $h$. Assume that $\pi$ is chosen to be the identity permutation. It is known [20] that the majority function possesses the best possible algebraic immunity. Motivated by this fact, we choose $h$ to be the majority function on $m$ variables. We prove that the resulting bent function on $2m$ variables has algebraic immunity at least $\lceil m/2\rceil$, and hence fast algebraic immunity at least $1+\lceil m/2\rceil$. On the implementation aspect, we show that the majority function can be computed using $O(m)$ NAND gates. So the obtained bent function has maximum nonlinearity, good algebraic immunity and at the same time is quite efficient to implement. The algebraic immunity is not the maximum possible, but due to the implementation efficiency, it is possible to increase the number of variables to achieve the desired level of algebraic resistance. One small problem is that a bent function is not balanced. This problem is easily rectified by XORing a new variable to the bent function, obtaining a function on an odd number of variables. This modification does not affect the linear bias or the algebraic resistance and requires only one extra XOR gate for implementation. A novelty of our work is the observation and the proof that choosing $h$ to be the majority function improves the algebraic immunity. While Maiorana-McFarland bent functions have been extensively studied in the literature, this simple observation had escaped the notice of earlier researchers.
@@
We perform a detailed concrete security analysis of some of the well known attacks on the nonlinear filter model. As the outcome of this analysis, for various security levels, we provide concrete proposals for stream ciphers based on the nonlinear filter model using the Boolean functions described above as the filtering functions. Further, we provide concrete gate count estimates for the entire circuit to implement the stream ciphers. For the 80-bit, $128$-bit, 160-bit, 192-bit, 224-bit and the $256$-bit security levels, we obtain gate count estimates of 1743.5, 2771.5, 3520.5, 4188.5, 4854.5, and 5607.5 NAND gates respectively. The gate count estimates for the 80-bit and the 128-bit security levels compare quite well [^1] Following [1] we estimated 8 NAND gates for a flip-flop, whereas Trivium estimated 12 NAND gates for a flip-flop. Using 12 NAND gates for a flip-flop, our proposal at the 80-bit security level requires about 2395.5 NAND gates, whereas Trivium requires about 3488 NAND gates. with famous ciphers such as Trivium [8] and Grain-128a [1] which offer 80-bit and 128-bit security respectively. For the other security levels, we are not aware of other stream ciphers which have such low gate counts. So our revival of the nonlinear filter model of stream ciphers leads to concrete proposals which offer a combination of both a desired level of security and low gate count.
@@
One advantage of using LFSRs is that it is possible to provably ensure that the LFSR has a maximum period. For stream ciphers based on NFSRs, such as Trivium and Mickey, such provable assurance is not available. For stream ciphers which use a combination of LFSR and NFSR such as Grain-128a, it is possible to mount a divide-and-conquer attack. For example the attack in [42] on Grain-128a finds the state of the LFSR indendently of the state of the NFSR. (In some ways this is reminiscent of the attack by Siegenthaler [41] on the nonlinear combiner model.) So even though Grain-128a uses a 256-bit state, due to the divide-and-conquer strategy the full protection of the large state is not achieved. On the other hand, for the nonlinear filter model, there is no known way to mount a divide-and-conquer attack. Our proposal at the 128-bit security level uses a 257-bit LFSR, and there is no way to estimate half the state of the LFSR without involving the other half.
@@ 2 Preliminaries
The fast algebraic immunity (FAI) of $f$ is a combined measure of resistance to both algebraic and fast algebraic attacks: ${\sf FAI}(f)\allowbreak=\allowbreak\min\left(2{\sf AI}(f),\allowbreak\min_{g\neq 0}\{{\sf deg}(g)\allowbreak+{\sf deg}(fg):\allowbreak 1\leq{\sf deg}(g)\allowbreak<\allowbreak{\sf AI}(f)\}\right)$. So ${\sf FAI}(f)\allowbreak=\allowbreak\min(\allowbreak 2{\sf AI}(f),\allowbreak\min\{e+d+1\}\allowbreak)$, where the second minimum is taken over all pairs $(e,d)$ in the FAA-profile of $f$. Further, for any function $f$, $1+{\sf AI}(f)\leq{\sf FAI}(f)\leq 2\,{\sf AI}(f)$.
@@ 3 Construction from Maiorana-McFarland Bent Functions
Suppose $m\geq 1$. There is an $\bm{\omega}^{\star}\in\mathbb{F}_{2}^{m}$ such that ${\sf AI}({\sf MM}_{2m})\geq{\sf wt}(\bm{\omega}^{\star})+{\sf AI}\big{(}\langle\bm{\omega}^{\star},\pi(\mathbf{X})\rangle\oplus h(\mathbf{X})\big{)}$.
Suppose that $\pi$ is a bit permutation, i.e. $\pi(X_{1},\ldots,X_{m})=(X_{\rho(1)},\ldots,X_{\rho(m)})$, for some permutation $\rho$ of $\{1,\ldots,m\}$. Then ${\sf AI}({\sf MM}_{2m})\geq{\sf AI}(h)$.
@@
By $({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m}$ we denote the function obtained by instantiating the definition of ${\sf MM}_{2m}$ given by (2) with $h={\sf Maj}_{m}$ and $\pi={\sf id}$. Similarly, by $({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m+1}$ we denote the $(2m+1)$-variable function obtained from $({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m}$ using (4).
@@ 4 Concrete Proposals
$f_{2m+1}(W,\mathbf{X},\mathbf{Y})$ $=$ $\left\{\begin{array}[]{ll}{\sf MM}_{2m+1}(W,\mathbf{X},\mathbf{Y})&\mbox{if }{\sf MM}_{2m+1}(\mathbf{0}_{2m+1})=1,\\ 1\oplus{\sf MM}_{2m+1}(W,\mathbf{X},\mathbf{Y})&\mbox{if }{\sf MM}_{2m+1}(\mathbf{0}_{2m+1})=0.\end{array}\right\}$ (7)
@@
The degree of $({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m+1}$ is $d=2^{\lfloor\log_{2}(2m+1)\rfloor}$. If $L$ is a prime, then the linear complexity of the generated keystream sequence is (${L\choose d}$ (see [37]) and so at least these many keystream bits are required to determine the linear complexity. So if
$\alpha={L\choose 2^{\lfloor\log_{2}(2m+1)\rfloor}}>2^{B},$ (8)
@@
Fast correlation attacks do not require exhaustive search on the states of the LFSR. There is a large literature on fast correlation attacks including older papers such as [35, 28, 29, 15, 12, 30, 11, 16] as well as more recent papers such as [42, 45, 44, 31, 32]. See [9, 10, 33, 2] for surveys of the area. In the following, we evaluate security against some representative fast correlation attacks and show how to choose the values of the design parameters so as to resist these attack. Recall from Proposition 2 that for ${\sf MM}_{2m+1}$ the linear bias $\varepsilon=2^{-m-1}$.
@@
Type of attack. The attacks in [30, 11] apply specifically to the filter generator model. These two attacks are essentially the same when the filtering function is plateaued (which is the case for $f_{2m+1}$). For the attack, $N$ is about $2^{(L-k)/w}$, and $T$ is about $2^{k}\cdot F^{w}$, where $w$ and $k$ are as in the attack in [15] and $F$ is the size of the support of the Walsh transform of the filtering function.
@@
Other fast correlation attacks. We next consider some of the more recent attacks. The attack in [42] is based on using $M>1$ linear approximations. For this attack, both $N$ and $T$ are about $2^{L-\beta}$, where $\beta$ is an algorithm parameter. A necessary condition for the attack to succeed is that $M>2^{\beta}$. For the attack to succeed at the $\kappa$-bit security level, i.e. $T\leq 2^{\kappa}$, it is required to have $\beta\geq L-\kappa$. From the bound on $M$, it follows that more than $2^{L-\kappa}$ linear approximations with sufficiently high correlations are required. For Grain-128a [1], $L$ equals 128 and about $2^{26.58}$ (i.e. $M$ is about $2^{26.58}$) linear approximations with absolute correlations at least $2^{-54.2381}$ were identified in [42]. For $\kappa=128$, Table 1 recommends $L=257$ and $m=58$. To apply the attack in [42] to $\mathcal{S}(257,59)$, more than $2^{129}$ linear approximations with sufficiently high correlations are required. The linear bias of the filtering function itself is $2^{-60}$. Finding multiple linear approximations requires combining keystream bits which further lowers the linear bias. So there is no approximation with linear bias greater than $2^{-60}$. Further, going through the details of the attack in [1], we could not identify any method to obtain more than $2^{129}$ linear approximations. So it is unlikely that the attack in [42] will apply to $\mathcal{S}(257,59)$.
Subsequent works (such as [45, 44, 31, 32]) on fast correlation attacks use vectorial decoding technique along with multiple linear approximations, use of the BKW algorithm [5], and multivariate correlation attack. The attacks are quite complex and there are no simple closed form expressions for the values of $N$ and time complexity. Upon studying the attacks in [42, 45, 44, 31, 32] The stream ciphers considered by these attacks are Grain-128a and Sosemanuk [4], which is another cipher providing 128-bit security. From the above discussion, we already know that the linear bias of $\mathcal{S}(257,59)$ is substantially lower than that of Grain-128a. For Sosemanuk, the best known [31] linear approximation has correlation $2^{-20.84}$, which is far greater than the linear bias $2^{-60}$ for $\mathcal{S}(257,59)$. The very low linear bias of $\mathcal{S}(257,59)$ and in general of the other stream cipher proposals in Table 1 make it quite unlikely that the attacks in [45, 44, 31, 32] will defeat the claimed security level of these stream ciphers.
@@
For a filtering function $f$ having ${\sf AI}(f)=a$, an algebraic attack requires about $N=\sum_{i=0}^{a}{L\choose i}$ keystream bits and has time complexity $T$ to be about $T^{\omega}$, where $\omega$ is the exponent of matrix multiplication (see Section 3.1.5 of [13]). We assume $\omega=2.8$. From Proposition 4, we conservatively assume that ${\sf AI}(({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m+1})=\lceil m/2\rceil$, i.e. the algebraic immunity is equal to its lower bound. Let
@@
For any $(e,d)$ in the FAA-profile of $({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m}$, the Berlekamp-Massey step in a fast algebraic attack on the filter generator model has time complexity $\mathcal{O}(ED\log D)$, where $E=\sum_{i=0}^{e}{L\choose i}$ and $D=\sum_{i=0}^{d}{L\choose i}$ (see [25] and Section 3.1.5 of [13]). This complexity dominates the overall time complexity of a fast algebraic attack and ignoring the logarithm term, and we take $T$ to be equal to $ED$. The maximum value of $e$ is one less than the algebraic immunity. The number $N$ of keystream bits required is about $2E$. From Proposition 4, we again assume that ${\sf AI}(({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m+1})=\lceil(2m+1)/4\rceil$. Recall that for any Boolean function, its fast algebraic immunity is at least one more than its algebraic immunity, and so ${\sf FAI}(({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m+1}))\geq 1+{\sf AI}(({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m+1})=1+\lceil(2m+1)/4\rceil$. For the concrete analysis, we use the lower bound $1+\lceil m/2\rceil$ as the actual value of ${\sf FAI}(({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m+1}))$ Let $a=\lceil m/2\rceil$, and define
$\gamma$ $=$ $\min_{\stackrel{{\scriptstyle 1\leq e\leq a\rfloor,}}{{d=a-e}}}\left(\sum_{i=0}^{e}{L\choose i}\right)\left(\sum_{i=0}^{d}{L\choose i}\right).$ (13)

### v2

@@ Abstract
The nonlinear filter model is an old and well understood approach to the design of secure stream ciphers. Extensive research over several decades has shown how to attack stream ciphers based on this model and has identified the security properties required of the Boolean function used as the filtering function to resist such attacks. This led to the problem of constructing Boolean functions which provide adequate security and at the same time are efficient to implement. Unfortunately, over the last two decades no good solutions to this problem appeared in the literature. The lack of good solutions has effectively led to nonlinear filter model becoming more or less obsolete. This is a big loss to the cryptographic design toolkit, since the great advantages of the nonlinear filter model are its simplicity, well understood security and the potential to provide low cost solutions for hardware oriented stream ciphers. In this paper we construct balanced functions on an odd number $n\geq 5$ of variables with the following provable properties: linear bias equal to $2^{-\lfloor n/2\rfloor-1}$, algebraic degree equal to $2^{\lfloor\log_{2}\lfloor n/2\rfloor\rfloor}$, algebraic immunity at least $\lceil(n-1)/4\rceil$, fast algebraic immunity at least $1+\lceil(n-1)/4\rceil$, and can be implemented using $O(n)$ NAND gates. The functions are obtained from a simple modification of the well known class of Maiorana-McFarland bent functions. By appropriately choosing $n$ and the length $L$ of the linear feedback shift register, we show that it is possible to obtain examples of stream ciphers which are provably $\kappa$-bit secure against well known classes of attacks for various values of $\kappa$. We provide concrete proposals for $\kappa=80,128,160,192,224$ and $256$. For the $80$-bit, $128$-bit, and the $256$-bit security levels, the circuits for the corresponding stream ciphers require about 1743.5, 2771.5, and 5607.5 NAND gates respectively. For the $80$-bit and the $128$-bit security levels, the gate count estimates compare quite well to the famous ciphers Trivium and Grain-128a respectively, while for the $256$-bit security level, we do not know of any other stream cipher design which has such a low gate count.
@@ 1 Introduction
Extensive research carried out over the last few decades has shown several approaches to cryptanalysing the nonlinear filter model of stream ciphers. The initial line of attack was based upon determining the linear complexity of the produced keystream. For an LFSR of length $L$ and a Boolean function of algebraic degree $d$, under certain reasonable and easy-to-ensure conditions, the linear complexity of the keystream is known to be at least (${L\choose d}$ (see [37]). Using large enough values of $L$ and $d$, linear complexity based attacks can be made infeasible. The second phase of attacks consisted of various kinds of (fast) correlation attacks. Starting with the first such attack in 1985 [41], a long line of papers [35, 23, 17, 28, 29, 15, 12, 30, 11, 16, 42, 45, 44, 31, 32] explored various avenues for mounting correlation attacks. Surveys of some of the older attacks appear in [9, 10, 33, 2]. The resistance to correlation attacks is mainly determined by the linear bias $\varepsilon$ of the Boolean function $f$. The linear bias is determined by the nonlinearity of the function $f$; the higher the nonlinearity, the lower the linear bias. The third phase of attacks started in 2003 with the publication of the algebraic attack [19] and was soon followed by the publication of the fast algebraic attack [18]. Resistance to these attacks requires the function $f$ to possess high (fast) algebraic immunity.
The various attacks mentioned above have posed the following design challenge for a Boolean function to be used in the nonlinear filter model of stream ciphers. Construct balanced Boolean functions which achieve a good combination of nonlinearity and algebraic resistance and are also very efficient to implement. Unfortunately, since the time algebraic attacks were proposed in the early-2000s, no good solutions to the design challenge for Boolean functions have appeared in the literature. A consequence of not being able to find good solutions to the design challenge is that the nonlinear filter model of stream ciphers became obsolete. This is somewhat unfortunate since the model is very old, well studied with well understood security, and the potential to provide low gate count solutions in hardware.
We note that even though the nonlinear filter model became obsolete, use of LFSRs in the design of stream cipher has continued for both hardware and software oriented proposals (see for example [26, 4, 22]). Instead of using a Boolean function, such designs typically use a nonlinear finite state machine to filter the output of the LFSR. Some ciphers such as [8, 3] have gone further and replaced the LFSR with one or more nonlinear feedback shift registers (NFSRs).
@@
The well known Maiorana-McFarland class of bent functions is defined as follows. For $m\geq 1$, let $\mathbf{X}$ and $\mathbf{Y}$ be two vectors of $m$ variables. Then a $2m$-variable Maiorana-McFarland bent function is defined to be $\langle\pi(\mathbf{X}),\mathbf{Y}\rangle\oplus h(\mathbf{X})$, where $\pi$ is a bijection from $m$-bit strings to $m$-bit strings and $h$ is any $m$-variable Boolean function. It is well-known that the nonlinearity of Maiorana-McFarland functions does not depend on the choice of $h$ nor on that of the permutation $\pi$. Bad choices of $\pi$ and $h$ (for example, $\pi$ to be the identity permutation and $h$ to be a constant function) provide functions whose algebraic immunity is low. On the bright side, we observe that it may be possible to improve the algebraic resistance by properly choosing $\pi$ and $h$. Assume that $\pi$ is chosen to be the identity permutation. It is known [20] that the majority function possesses the best possible algebraic immunity. Motivated by this fact, we choose $h$ to be the majority function on $m$ variables. We prove that the resulting bent function on $2m$ variables has algebraic immunity at least $\lceil m/2\rceil$, and hence fast algebraic immunity at least $1+\lceil m/2\rceil$. On the implementation aspect, we show that the majority function can be computed using $O(m)$ NAND gates. So the obtained bent function has maximum nonlinearity, good algebraic immunity and at the same time is quite efficient to implement. The algebraic immunity is not the maximum possible, but due to the implementation efficiency, it is possible to increase the number of variables to achieve the desired level of algebraic resistance. One small problem is that a bent function is not balanced. This problem is easily rectified by XORing a new variable to the bent function, obtaining a function on an odd number of variables. This modification does not affect the linear bias or the algebraic resistance and requires only one extra XOR gate for implementation. A novelty of our work is the observation and the proof that choosing $h$ to be the majority function improves the algebraic immunity. While Maiorana-McFarland bent functions have been extensively studied in the literature, this simple observation has escaped the notice of earlier researchers.
@@
We perform a detailed concrete security analysis of some of the well known attacks on the nonlinear filter model. As the outcome of this analysis, for various security levels, we provide concrete proposals for stream ciphers based on the nonlinear filter model using the Boolean functions described above as the filtering functions. A strong point in favour of these proposals is that at the appropriate security levels they provide provable assurance against well known classes of attacks. Further, we provide concrete gate count estimates for the entire circuit to implement the stream ciphers. For the 80-bit, $128$-bit, 160-bit, 192-bit, 224-bit and the $256$-bit security levels, we obtain gate count estimates of 1743.5, 2771.5, 3520.5, 4188.5, 4854.5, and 5607.5 NAND gates respectively. The gate count estimates for the 80-bit and the 128-bit security levels compare quite well [^1] Following [1] we estimated 8 NAND gates for a flip-flop, whereas Trivium estimated 12 NAND gates for a flip-flop. Using 12 NAND gates for a flip-flop, our proposal at the 80-bit security level requires about 2395.5 NAND gates, whereas Trivium requires about 3488 NAND gates. with famous ciphers such as Trivium [8] and Grain-128a [1] which offer 80-bit and 128-bit security respectively. For the other security levels, we are not aware of other stream ciphers which have such low gate counts. So our revival of the nonlinear filter model of stream ciphers leads to concrete proposals which offer a combination of both provable security against well known classes of attacks at a desired level of security and also low gate count.
@@
One advantage of using LFSRs is that it is possible to provably ensure that the LFSR has a maximum period. For stream ciphers based on NFSRs, such as Trivium and Mickey, such provable assurance is not available. For stream ciphers which use a combination of LFSR and NFSR such as Grain-128a, it is possible to mount a divide-and-conquer attack. For example the attack in [42] on Grain-128a finds the state of the LFSR independently of the state of the NFSR. (In some ways this is reminiscent of the attack by Siegenthaler [41] on the nonlinear combiner model.) So even though Grain-128a uses a 256-bit state, due to the divide-and-conquer strategy the full protection of the large state is not achieved. On the other hand, for the nonlinear filter model, there is no known way to mount a divide-and-conquer attack. Our proposal at the 128-bit security level uses a 257-bit LFSR, and there is no way to estimate half the state of the LFSR without involving the other half.
@@ 2 Preliminaries
The fast algebraic immunity (FAI) of $f$ is a combined measure of resistance to both algebraic and fast algebraic attacks: ${\sf FAI}(f)\allowbreak=\allowbreak\min\left(2{\sf AI}(f),\allowbreak\min_{g\neq 0}\allowbreak\{{\sf deg}(g)\allowbreak+\allowbreak{\sf deg}(fg):\allowbreak 1\allowbreak\leq\allowbreak{\sf deg}(g)\allowbreak<\allowbreak{\sf AI}(f)\}\right)$.
So ${\sf FAI}(f)\allowbreak=\allowbreak\min(\allowbreak 2{\sf AI}(f),\allowbreak\min\{e+d+1\}\allowbreak)$, where the second minimum is taken over all pairs $(e,d)$ in the FAA-profile of $f$. Further, for any function $f$, $1+{\sf AI}(f)\leq{\sf FAI}(f)\leq 2\,{\sf AI}(f)$.
@@ 3 Construction from Maiorana-McFarland Bent Functions
Suppose $m\geq 1$. There is an $\bm{\omega}^{\star}\in\mathbb{F}_{2}^{m}$ such that
${\sf AI}({\sf MM}_{2m})$$\geq$ ${\sf wt}(\bm{\omega}^{\star})+{\sf AI}\big{(}\langle\bm{\omega}^{\star},\pi(\mathbf{X})\rangle\oplus h(\mathbf{X})\big{)}.$
Suppose that $\pi$ is a bit permutation, i.e. $\pi(X_{1},\ldots,X_{m})=(X_{\rho(1)},\ldots,X_{\rho(m)})$, for some permutation $\rho$ of $\{1,\ldots,m\}$. Then
${\sf AI}({\sf MM}_{2m})$$\geq$ ${\sf AI}(h).$
@@
By $({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m}$ we denote the function obtained by instantiating the definition of ${\sf MM}_{2m}$ given by (2) with $h={\sf Maj}_{m}$ and $\pi={\sf id}$. Similarly, by $({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m+1}$, we denote the function obtained from $({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m}$ using (4).
@@ 4 Concrete Proposals
$f_{2m+1}(W,\mathbf{X},\mathbf{Y})$ $=$ $\left\{\begin{array}[]{ll}{\sf MM}_{2m+1}(W,\mathbf{X},\mathbf{Y})&\mbox{if }{\sf MM}_{2m+1}(\mathbf{0}_{2m+1})=1,\\ 1\oplus{\sf MM}_{2m+1}(W,\mathbf{X},\mathbf{Y})&\mbox{if }{\sf MM}_{2m+1}(\mathbf{0}_{2m+1})=0.\end{array}\right.$ (7)
@@
The degree of $({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m+1}$ is $d=2^{\lfloor\log_{2}m\rfloor}$. If $L$ is a prime, then the linear complexity of the generated keystream sequence is (${L\choose d}$ (see [37]) and so at least these many keystream bits are required to determine the linear complexity. So if
$\alpha={L\choose 2^{\lfloor\log_{2}m\rfloor}}>2^{B},$ (8)
@@
Fast correlation attacks do not require exhaustive search on the states of the LFSR. There is a large literature on fast correlation attacks including older papers such as [35, 28, 29, 15, 12, 30, 11, 16] as well as more recent papers such as [42, 45, 44, 31, 32]. See [9, 10, 33, 2] for surveys of the area. In the following, we evaluate security against some representative fast correlation attacks and show how to choose the values of the design parameters so as to resist these attack. Recall from Proposition 2 that for ${\sf MM}_{2m+1}$, and hence for $f_{2m+1}$, the linear bias $\varepsilon=2^{-m-1}$.
@@
Type of attack. The attacks in [30, 11] apply specifically to the filter generator model. These two attacks are essentially the same when the filtering function is plateaued (which is the case for $f_{2m+1}$). For the attack, $N$ is about $2^{(L-k)/w}$, and $T$ is about $2^{k}\cdot F^{w}$, where $w$ and $k$ are as in the attack in [15] (see above) and $F$ is the size of the support of the Walsh transform of the filtering function.
@@
Other fast correlation attacks. We next consider some of the more recent attacks. The attack in [42] is based on using $M>1$ linear approximations. For this attack, both $N$ and $T$ are about $2^{L-\beta}$, where $\beta$ is an algorithm parameter. A necessary condition for the attack to succeed is that $M>2^{\beta}$. For the attack to succeed at the $\kappa$-bit security level, i.e. $T\leq 2^{\kappa}$, it is required to have $\beta\geq L-\kappa$. From the bound on $M$, it follows that more than $2^{L-\kappa}$ linear approximations with sufficiently high correlations are required. For Grain-128a [1], $L$ equals 128 and about $2^{26.58}$ (i.e. $M$ is about $2^{26.58}$) linear approximations with absolute correlations at least $2^{-54.2381}$ were identified in [42]. For $\kappa=128$, Table 1 recommends $L=257$ and $m=58$. To apply the attack in [42] to $\mathcal{S}(257,59)$, more than $2^{129}$ linear approximations with sufficiently high correlations are required. The linear bias of the filtering function itself is $2^{-60}$. Finding multiple linear approximations requires combining keystream bits which further lowers the linear bias. So there is no approximation with linear bias greater than $2^{-60}$. Further, going through the details of the attack in [1], we could not identify any method to obtain more than $2^{129}$ linear approximations. So there does not seem to be any way to apply the attack in [42] to $\mathcal{S}(257,59)$.
Subsequent works (such as [45, 44, 31, 32]) on fast correlation attacks use vectorial decoding technique along with multiple linear approximations, use of the BKW algorithm [5], and multivariate correlation attack. The attacks are quite complex and there are no simple closed form expressions for the values of $N$ and $T$. The stream ciphers to which these attacks are applied are Grain-128a and Sosemanuk [4], which is another cipher providing 128-bit security. From the above discussion, we already know that the linear bias of $\mathcal{S}(257,59)$ is substantially lower than that of Grain-128a. For Sosemanuk, the best known [31] linear approximation has correlation $2^{-20.84}$, which is far greater than the linear bias $2^{-60}$ for $\mathcal{S}(257,59)$. The very low linear bias of $\mathcal{S}(257,59)$ and in general of the other stream cipher proposals in Table 1 make the attacks in [45, 44, 31, 32] inapplicable at the stated security levels.
@@
For a filtering function $f$ having ${\sf AI}(f)=a$, an algebraic attack requires about $N=\sum_{i=0}^{a}{L\choose i}$ keystream bits and has time complexity $T$ to be about $N^{\omega}$, where $\omega$ is the exponent of matrix multiplication (see Section 3.1.5 of [13]). We assume $\omega=2.8$. From Proposition 4, we conservatively assume that ${\sf AI}(({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m+1})=\lceil m/2\rceil$, i.e. the algebraic immunity is equal to its lower bound. Let
@@
For any $(e,d)$ in the FAA-profile of $({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m}$, the Berlekamp-Massey step in a fast algebraic attack on the filter generator model has time complexity $\mathcal{O}(ED\log D)$, where $E=\sum_{i=0}^{e}{L\choose i}$ and $D=\sum_{i=0}^{d}{L\choose i}$ (see [25] and Section 3.1.5 of [13]). This complexity dominates the overall time complexity of a fast algebraic attack and ignoring the logarithm term, and we take $T$ to be equal to $ED$. The maximum value of $e$ is one less than the algebraic immunity. The number $N$ of keystream bits required is about $2E$. From Proposition 4, we again assume that ${\sf AI}(({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m+1})=\lceil(2m+1)/4\rceil$. Recall that for any Boolean function, its fast algebraic immunity is at least one more than its algebraic immunity, and so ${\sf FAI}(({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m+1}))\geq 1+{\sf AI}(({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m+1})=1+\lceil(2m+1)/4\rceil$. For the concrete analysis, we use the lower bound $1+\lceil m/2\rceil$ as the actual value of ${\sf FAI}(({\sf Maj},{\sf id})\mbox{-}{\sf MM}_{2m+1}))$ Define
$\gamma$ $=$ $\min_{\stackrel{{\scriptstyle 1\leq e\leq\lceil m/2\rceil\rfloor,}}{{d=a-e}}}\left(\sum_{i=0}^{e}{L\choose i}\right)\left(\sum_{i=0}^{d}{L\choose i}\right).$ (13)